When a ball is thrown, its height in feet h after t seconds is given by the equation
h=vt-16t^2

where v is the initial upwards velocity in feet per second. If v=15 feet per second, find all values t of for which h=3 feet. Do not round any intermediate steps. Round your answer to 2 decimal places.

Answer by stanbon(75887)

When a ball is thrown, its height in feet h after t seconds is given by the equation
h=vt-16t^2
where v is the initial upwards velocity in feet per second. If v=15 feet per second, find all values t of for which h=3 feet.
----------------
h(t) = vt-16t^2
3 = 15t - 16t^2
16t^2 - 15t + 3 = 0
----
Use the Quadratic formula to solve for "t":
t = [15 +- sqrt(225 - 4*16*3)]/32
t = [15 +- sqrt(33)]/32
t = [15 + sqrt(33)]/32 = 0.64827 seconds
or
t = [15 - sqrt(33)]/32 = 0.28923 seconds
================================
Cheers,
Stan H.